Solve the following equation by completing the square.
9x^2-6x=2
The solution set is ___

Answer by jorel1380(3719)

9x²-6x=2
9x²-6x+1=2+1
(3x-1)²=3
3x-1=√3
3x=1+/-√3
x=(1+/-√3)/3. ☺☺☺☺